OpenStudy (mislovely187):

Find the length of the radius of the given circle. (x - 5)2 + (y + 3)2 = 25

OpenStudy (anonymous):

you have (x−h) 2 +(y−k) 2 =r 2 the equation for a circle with radius r since r 2 =25 you know r=5

OpenStudy (mislovely187):

thank you

OpenStudy (anonymous):

your welcome

OpenStudy (victoriasushchik):

All you need to do is complete two squares. x2 + y2 - 4x - 6y - 12 = 0 x2 - 4x + y2 - 6y = 12 x2 - 4x + 4 + y2 - 6y + 9 = 12 + 4 + 9 (x - 2)2 + (y - 3)2 = 25 = 52 Thus it has the center at (2,3) and a radius r = 5.